# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ C.1 Conceptual Framework
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is households consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
@@
Subject to equation(20), the social planner maximizes equation(19) by choosing an allocation of consumption across households. The suggested first-order condition at time $t$ is derived as follows:
@@
where $\lambda$ is the Lagrange multiplier for the single lifetime resource constraint.
To derive my empirical specification, I assume a constant relative risk aversion preference as:
@@
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (e.g., Hayashi et al. 1996). This simplifies the notation in equation 21 as:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income. [^92] In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$. The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
Given this technical issue, I used the adjusted monthly panel dataset to analyze the estimate for the rent category. I confirmed that the adjusted monthly panel dataset (Online Appendix B.7) can assign the rent payday each month. The scatter plot from the adjusted monthly variation shows a flat relationship in Figure C.4c.
Figure C.2: Raw Relationship between the Log-differences in Semi-monthly Income and Expenditure
@@
Figure C.2: Raw Relationship between the Log-differences in Semi-monthly Income and Expenditure
@@
Figure C.4: Raw Relationship between the Log-differences in Monthly Income and Expenditure
@@
Figure C.4: Raw Relationship between the Log-differences in Monthly Income and Expenditure
@@
  • [1] Arellano, Manuel. “Computing standard errors for robust within-groups estimators.” Oxford Bulletin of Economics and Statistics 49, no. 4 (1987): 431–434.
@@
  • [2] Cochrane, John H. “A Simple Test of Consumption Insurance.” Journal of Political Economy 99, no. 5 (1991): 957–976.
@@
  • [3] Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
@@
  • [4] James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
@@
  • [5] Jappelli, Tullio, and Luigi Pistaferri. The economics of consumption: Theory and evidence. New York, NY: Oxford University Press, 2017.
@@
  • [6] Kawai, Takao. “Shakai chōsa houhoushi ni tsuite (On history of social research methods)” [in Japanese]. Hōgaku Kenkyū 53, no. 9 (1980): 41–67.
@@
  • [7] Kawai, Takao. “Tsukishima chōsa saikōsatsu, ichi (Tsukishima Survey reconsidered (1))” [in Japanese]. Hōgaku Kenkyū 54, no. 8 (1981a): 1–26.
@@
  • [8] Kawai, Takao. “Tsukishima chōsa saikōsatsu, ni kan (Tsukishima Survey reconsidered (2, end))” [in Japanese]. Hōgaku Kenkyū 54, no. 9 (1981b): 1–33.
@@
  • [9] Mace, Barbara J. “Full Insurance in the Presence of Aggregate Uncertainty.” Journal of Political Economy 99, no. 5 (1991): 928–956.
@@
  • [10] Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences.” American Economic Review 102, no. 1 (2012): 428–468.
@@
  • [11] Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Reply.” American Economic Review 108, no. 10 (2018): 3114–3115.
@@
  • [12] Miyoshi, Toyotaro. “Tsukishima chōsa no seiritsu to sonokeika ni tsuite (The establishment and progress of the Tsukishima Survey)” [in Japanese]. Meiseidaigaku Kenkyū Kiyō 16, (1980): 31–48.
@@
  • [13] Takano, Iwasaburo. Honpō shakaitōkei ron (Social Statistics in Japan) [in Japanese]. Tokyo: Kaizō sha.
@@
  • [14] Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
@@
  • [15] Sekiya, Koichi. Seikatsu koten sōsho, volume 6: Tsukishima chōsa (A classic series of books about life, volume 6: Tsukishima Survey) [in Japanese]. Tokyo: Kōseikan.
@@
  • [16] Shrinivas, Aditya and Marcel Fafchamps. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Comment.” American Economic Review 108, no. 10 (2018): 3104–3113.
@@
  • [17] Shibuya, Ryuichi, Kameji Suzuki, and Shojiro Ishiyama. Pawn Shops in Japan. [in Japanese] Tokyo: Waseda University Press, 1982.
@@
  • [18] Shibuya, Ryuichi. Development and Structure of Usuries. [in Japanese] Tokyo: Nihon Tosho Center, 2000.
@@
  • [19] Takeda, Naoko. “Shitsuteki chōsa data no nijibunseki (Qualitative Secondary Analysis and Reuse of the Original Research Data Set of the Japanese Sociological Classic)” [in Japanese]. Nihonrōdōkenkyūzasshi 57, no. 12 (2015): 70–80.
@@
  • [20] Tanaka, Hikaru. Another Financial System [in Japanese]. Nagoya: Nagoya University Press, 2018.
@@
  • [21] Terade, Kouji. “Tsukishima hōkokusho dai nishū (A study of “The volume II of the report on the investigation at Tsukishima”)” [in Japanese]. Mitagakkaizasshi 75, no. 6 (1982): 94–108.
@@
  • [22] Odaka, Konosuke. “Japanese-style Labour Relations.” In The Japanese Economic System and Its Historical Origins, edited by Tetsuji Okazaki and Masahiro Okuno-Fujiwara. Oxford: Oxford University Press, 1999.
@@
  • [23] Patrick, Royston and Douglas G. Altman. “Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
@@
  • [24] Wooldridge, Jeffrey M. Econometric Analysis of Cross Section and Panel Data. Cambridge: MIT Press, 2002.
@@
  • [1] Central Federation of Industrial Associations. Sangyō kumiai chōsa siryō, volume 9 (Survey report of the industrial association, volume.9). [in Japanese] Tokyo: Central Federation of Industrial Associations, 1925.
@@
  • [2] Committee for the History of the Labor Movement. Nihon rōdōundō shiryō (Historical Documents of the Japanese Labor Movement, volume 10). [in Japanese] Tokyo: Committee for the History of the Labor Movement, 1964.
@@
  • [3] Chūō Ward. Chūō ku enkaku zushū, Tsukishima hen (Chūō Ward History Chart, Tsukishima edition). [in Japanese] Tokyo: Jinbun sha, 1994.
@@
  • [4] Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923a.
@@
  • [5] Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ichi) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 1 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923b.
@@
  • [6] Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ni) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 2 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923c.
@@
  • [7] Ito, Osamau, and Saito Nao. Kinyu gyo(Financial Industry) [in Japanese]. Tokyo: Nihon Keieishi Kenkyūjyo, 2019.
@@
  • [8] Japan Meteorological Agency (database) [in Japanese]. https://www.data.jma.go.jp/gmd/risk/obsdl/index.php.
@@
  • [9] Kitazawa, Shinjiro. Tokyo niokeru kikaikogyō no jyukurenshokkō toshiteno shiagekō narabini senbankō no chinginhōkoku (The report on a survey of wages of fitters and turners as skilled-workers in the machine industry in Tokyo). [in Japanese] Tokyo: Ohara Institute of Social Research, 1924.
@@
  • [10] Kyoto City Office.Kyōto shi tōkeisho, dai jyūsan kai (Statistical Table of Kyoto City, volume. 13). [in Japanese] Kyoto: Kyoto City Office, 1922.
@@
  • [11] Kyoto Prefecture.Kyōto fu tōkeisho, taishō kyū nen, dai san hen (Statistical Table of Kyoto Prefecture, 1920 edition, vol. 3). [in Japanese] Kyoto: Kyoto Prefecture, 1922.
@@
  • [12] Ministry of Agriculture and Commerce. Kōjyō Tsūran (Taishō jyūnen jyūgatsu) (Handbook of Factories, November 1921 edition) [in Japanese]. Tokyo: Ministry of Agriculture and Commerce, 1921.
@@
  • [13] Ministry of Posts and Telecommunications. Denki jigyō yōran, dai jyūyon kai (Handbook of Electricity Industry, vol. 14). [in Japanese] Tokyo: Electric Association, 1922.
@@
  • [14] Nagoya City Office.Nagoya shi tōkeisho, dai nijyūni kai (Statistical Table of Kyoto City, volume. 22). [in Japanese] Nagoya: Nagoya City Office, 1922.
@@
  • [15] Ohkawa, Kazushi, Miyohei Shinohara, and Mataji Umemura. Estimates of Long-term Economic Statistics of Japan since 1868, volume.8. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1967.
@@
  • [16] Osaka City Office. Osakashi tōkeisho, Taishō kyūnen (Statistical Table of Osaka City, 1920 Edition). [in Japanese] Osaka: Osaka City Office, 1922.
@@
  • [17] Postal Savings Bureau. Chochikukyoku nenpō, volume 30 (Annual Statistical Report of the Postal Savings Bureau, volume 30). [in Japanese] Postal Savings Bureau, 1924.
@@
  • [18] Statistics Bureau of the Cabinet. Taishō kyūnen kokusei chōsa hōkoku, zenkoku no bu, dai ikkan (Population Census of Japan, Vol. 1, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1928.
@@
  • [19] Statistics Bureau of the Cabinet. Taoshō kyūnen kokuseichōsahōkoku, zenkoku no bu, dai nikan (Population Census of Japan, Vol. 2, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1929a.

### 25 Feb 2024 v6

@@ C.1 Conceptual Framework
where $\omega_{i}$ is the social planner’s weight representing the reciprocal of the marginal utility of each household ($10<\omega_{i}<$ 1 and $\sum_{i=1}^{N}\omega_{i}=1$), $10<\rho^{t}<$ 1 is the discount factor, $c_{i,t}(s_{t})$ is consumption, and $\theta_{i,t}(s_{t})$ is a preference shock. For all states $s_{t}$, the aggregate resource constraint is given by:
@@
Following Mace (1991), I assume that the aggregate idiosyncratic shocks approach zero as $N$ gets large. Consider the situation that the household’s endowment can be decomposed into the permanent component ($e_{i,t}^{P}$), shock linked to macroeconomic trends ($\varrho_{i,t}^{a}(s_{t})$), and idiosyncratic shock ($\varsigma_{i,t}(s_{t})$) as:
$e_{i,t}(s_{t})=e_{i,t}^{P}+\varrho_{i,t}^{a}(s_{t})+\varsigma_{i,t}(s_{t}).$ (22)
@@
Subject to equation(21), the social planner maximizes equation(20) by choosing an allocation of consumption across households. The first-order condition at time $t$ is derived as follows:
@@
where $\lambda$ is the Lagrange multiplier for the single lifetime resource constraint. To derive my empirical specification, I assume a constant relative risk aversion preference as:
@@
where $\sigma>0$ is the coefficient of relative risk aversion. Below, I consider a particular history of the state of the world so that the subscript $t$ can represent the dependence on the state (Hayashi et al. 1996). This simplifies the notation in equation 23 as:
@@
where $\tilde{y}_{i,t}$ is head’s earning, $x_{i,t}$ is the family size control, $\nu_{i}$ indicates the household-fixed effect, $\zeta_{t}$ indicates the time-fixed effect, and $\epsilon_{i,t}$ is a random error term. I herein use the head’s earnings ($\tilde{y}_{i,t}$) as the idiosyncratic shock variable instead of the disposable income.100100100In the analysis, I consider the labor-supply adjustment variable and sales of miscellaneous assets as the additional dependent variables in addition to the net income variables $r_{i,t}$ (Section 4.2). The former variable is the total earnings of the family members except for the head, meaning that the simultaneous issue rises if I use the disposable income in this equation. The regression is separately run for each component in the net income vector ($\bm{z}$).
@@ C.2 Raw Relationship between Income and Expenditures
Given this technical issue, the adjusted monthly panel dataset is suitable for estimating the income elasticity for the rent category. In fact, Figure C.4c confirms that the scatter plot from the adjusted monthly variation shows a flatter relationship.
@@
Figure C.2: Raw Relationship between the Log-differences
in Semi-monthly Income and Expenditure
@@
Figure C.2: Raw Relationship between the Log-differences
in Semi-monthly Income and Expenditure
@@
Figure C.4: Raw Relationship between the Log-differences
in Monthly Income and Expenditure
@@
Figure C.4: Raw Relationship between the Log-differences
in Monthly Income and Expenditure
C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
Table C.1 summarizes the contribution of regular customers to their total sales, social classes of the customers, and sales methods in the retailers selling white rice, fish, greengrocers, and fuels (firewood and charcoal) by scale of business. In the original document, the figures for “Total” in the firewood and charcoal category are missing in “Sales Method (%)”. For these three cells, I filled the averages over the three business-size categories in each sales method: 20.4% for “Cash”, 78.4% for “Credit purchase”, and 1.3% for “Monthly installment”.
Table C.1: Customers, Sales Methods, and Competitiveness in the Retailers in Tokyo City
Selling Item Scale of Number of Share of sales Social Class of Sales Method (%) Number of
Business stores for regular the Regular Customer (%) Cash Credit Monthly competitors
reported customer (%) Upper Middle Other purchase installment reported
White Rice Small 19 69.7 12.1 49.5 38.4 29.5 70.2 0.3 7.3
Medium 55 69.5 17.5 41.6 40.9 26.9 72.6 0.5 8.7
Large 27 81.7 15.9 51.5 32.6 26.7 71.0 2.3 8.8
Total 101 73.1 16 45.7 38.2 26.9 71.5 1.7 8.5
Fish Small 34 64.5 5.9 53.2 40.9 42.7 57.3 0 5.1
Medium 20 62.6 10 58.0 32.0 43.8 56.2 0 5.3
Large 5 72.0 0 86.0 14.0 44.8 55.2 0 3.4
Total 59 64.6 6.8 57.6 35.6 43.9 56.1 0 5.0
Greengrocer Small 46 42.2 8 39.2 52.7 59.7 40.3 0 6.5
Medium 17 57.1 8.8 43.5 47.6 65.0 35.0 0 6.1
Large 3 36.7 33.3 40.0 26.7 79.7 20.3 0 3.7
Total 66 45.4 9.4 40.4 50.2 65.8 34.2 0 6.2
Firewood and charcoal Small 78 68.5 13.5 50.7 35.8 25.0 74.2 0.8 8.4
Medium 17 77.7 23.5 49.1 27.4 19.3 79.1 1.6 8.9
Large 5 33.3 23.0 60.0 17.0 16.8 81.8 1.4 10.6
Total 100 51.1 15.7 50.9 33.4 20.4 78.4 1.3 8.6
Notes: This table summarizes the statistics on the sales in the retail shops selling white rice, fish, fruit and vegetables, and firewood and charcoal in Tokyo surveyed in 1935. 1. Small, middle, and large-scale businesses include retailers with annual sales of less than 10,000, 10,000–30,000, and 30,000+ yen, respectively. 2. The number of stores reported the share of regular customers (%) are 90, 56, 61, and 93 in the white rice, fish, greengrocers, and firewood and charcoal categories, respectively. 3. The share of regular customers is defined as the percentage of sales for regular customers in the annual total sales in 1935. 4. The definition of the social classes (“Upper”; “Middle”; “Other”) is not clearly defined in the report. Therefore, this must be based on the subjective judgment of the retailers. 5. The share of sales method is based on the percentage of sales using each method in the annual total sales in 1935. 6. The number of competitors indicates the number of peers within a three-cho (approximately 327 meters) radius. Source: Tokyo City Chamber of Commerce and Industry 1937, pp. legend; 20–21; questionnaire.
@@ C.3 Retail Sales in Tokyo City: A Retail Store Survey of 1935
  • (2) Arellano, Manuel. “Computing standard errors for robust within-groups estimators.” Oxford Bulletin of Economics and Statistics 49, no. 4 (1987): 431–434.
  • (3)
@@
  • (4) Cochrane, John H. “A Simple Test of Consumption Insurance.” Journal of Political Economy 99, no. 5 (1991): 957–976.
  • (5)
@@
  • (6) Choi, In “Unit root tests for panel data”. Journal of International Money and Finance 20, no. 2 (2001): 249–272.
  • (7)
@@
  • (8) James, John A., and Isao Suto. “Early Twentieth-century Japanese Worker Saving: Precautionary Behaviour before a Social Safety Net.” Cliometrica 5, no. 1 (2011): 1–25.
  • (9)
@@
  • (10) Jappelli, Tullio, and Luigi Pistaferri. The economics of consumption: Theory and evidence. New York, NY: Oxford University Press, 2017.
  • (11)
@@
  • (12) Kawai, Takao. “Shakai chōsa houhoushi ni tsuite (On history of social research methods)” [in Japanese]. Hōgaku Kenkyū 53, no. 9 (1980): 41–67.
  • (13)
@@
  • (14) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ichi (Tsukishima Survey reconsidered (1))” [in Japanese]. Hōgaku Kenkyū 54, no. 8 (1981a): 1–26.
  • (15)
@@
  • (16) Kawai, Takao. “Tsukishima chōsa saikōsatsu, ni kan (Tsukishima Survey reconsidered (2, end))” [in Japanese]. Hōgaku Kenkyū 54, no. 9 (1981b): 1–33.
  • (17)
@@
  • (18) Mace, Barbara J. “Full Insurance in the Presence of Aggregate Uncertainty.” Journal of Political Economy 99, no. 5 (1991): 928–956.
  • (19)
@@
  • (20) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences.” The American Economic Review 102, no. 1 (2012): 428–468.
  • (21)
@@
  • (22) Mazzocco, Maurizio and Shiv Saini. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Reply.” The American Economic Review 108, no. 10 (2018): 3114–3115.
  • (23)
@@
  • (24) Miyoshi, Toyotaro. “Tsukishima chōsa no seiritsu to sonokeika ni tsuite (The establishment and progress of the Tsukishima Survey)” [in Japanese]. Meiseidaigaku Kenkyū Kiyō 16, (1980): 31–48.
  • (25)
@@
  • (26) Takano, Iwasaburo. Honpō shakaitōkei ron (Social Statistics in Japan) [in Japanese]. Tokyo: Kaizō sha.
  • (27)
  • (28) Pearson, Karl. On the criterion that a given system of deviations from the probable in the case of a correlated system of variables is such that it can be reasonably supposed to have arisen from random sampling. Philosophical Magazine 50, no. 302, (1900): 157–175.
  • (29)
@@
  • (30) Schulhofer-Wohl, Sam. “Heterogeneity and Tests of Risk Preference”. Journal of Political Economy 119, no.5, (2011): 925–958.
  • (31)
@@
  • (32) Sekiya, Koichi. Seikatsu koten sōsho, volume 6: Tsukishima chōsa (A classic series of books about life, volume 6: Tsukishima Survey) [in Japanese]. Tokyo: Kōseikan.
  • (33)
@@
  • (34) Shrinivas, Aditya and Marcel Fafchamps. “Testing Efficient Risk Sharing with Heterogeneous Risk Preferences: Comment.” The American Economic Review 108, no. 10 (2018): 3104–3113.
  • (35)
@@
  • (36) Shibuya, Ryuichi, Kameji Suzuki, and Shojiro Ishiyama. Nihon no Shichiya (Pawn Shops in Japan) [in Japanese]. Tokyo: Waseda University Press, 1982.
  • (37)
@@
  • (38) Shibuya, Ryuichi. Kourigashikinyū no tenkai kōzō (Development and Structure of Usuries) [in Japanese]. Tokyo: Nihon Tosho Center, 2000.
  • (39)
@@
  • (40) Takeda, Naoko. “Shitsuteki chōsa data no nijibunseki (Qualitative Secondary Analysis and Reuse of the Original Research Data Set of the Japanese Sociological Classic)” [in Japanese]. Nihonrōdōkenkyūzasshi 57, no. 12 (2015): 70–80.
  • (41)
@@
  • (42) Tanaka, Hikaru. Mouhitotsu no kinyū system (Another Financial System) [in Japanese]. Nagoya: Nagoya University Press, 2018.
  • (43)
@@
  • (44) Terade, Kouji. “Tsukishima hōkokusho dai nishū (A study of “The volume II of the report on the investigation at Tsukishima”)” [in Japanese]. Mitagakkaizasshi 75, no. 6 (1982): 94–108.
  • (45)
@@
  • (46) Odaka, Konosuke. “Japanese-style Labour Relations.” In The Japanese Economic System and Its Historical Origins, edited by Tetsuji Okazaki and Masahiro Okuno-Fujiwara. Oxford: Oxford University Press, 1999.
  • (47)
@@
  • (48) Patrick, Royston and Douglas G. Altman. “Regression using Fractional Polynomials of Continuous Covariates: Parsimonious Parametric Modeling.” Journal of the Royal Statistical Society. Series C (Applied Statistics) 43 no. 3 (1994): 429–467.
  • (49)
@@
  • (50) Wooldridge, Jeffrey M. Econometric Analysis of Cross Section and Panel Data. Cambridge: MIT Press, 2002.
  • (51)
@@
  • (4) Central Federation of Industrial Associations. Sangyō kumiai chōsa siryō, volume 9 (Survey report of the industrial association, volume.9). [in Japanese] Tokyo: Central Federation of Industrial Associations, 1925.
  • (5)
@@
  • (6) Committee for the History of the Labor Movement. Nihon rōdōundō shiryō (Historical Documents of the Japanese Labor Movement, volume 10). [in Japanese] Tokyo: Committee for the History of the Labor Movement, 1964.
  • (7)
@@
  • (8) Chūō Ward. Chūō ku enkaku zushū, Tsukishima hen (Chūō Ward History Chart, Tsukishima edition). [in Japanese] Tokyo: Jinbun sha, 1994.
  • (9)
@@
  • (10) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923a.
  • (11)
@@
  • (12) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ichi) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 1 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923b.
  • (13)
@@
  • (14) Department of Health, Ministry of the Interior. Tokyo-shi Kyobashi-ku Tsukishima ni okeru jicchichōsahōkoku (dai isshū, furoku no ni) (Report of the Field Study in Tsukishima in Kyobashi Ward of Tokyo City, Appendix 2 of Volume 1) [in Japanese]. Tokyo: Department of Health, Ministry of the Interior, 1923c.
  • (15)
@@
  • (16) Ito, Osamau, and Saito Nao. Kinyu gyo(Financial Industry) [in Japanese]. Tokyo: Nihon Keieishi Kenkyūjyo, 2019.
  • (17)
@@
  • (18) Japan Meteorological Agency (database) [in Japanese]. https://www.data.jma.go.jp/gmd/risk/obsdl/index.php.
  • (19)
@@
  • (20) Kitazawa, Shinjiro. Tokyo niokeru kikaikogyō no jyukurenshokkō toshiteno shiagekō narabini senbankō no chinginhōkoku (The report on a survey of wages of fitters and turners as skilled-workers in the machine industry in Tokyo). [in Japanese] Tokyo: Ohara Institute of Social Research, 1924.
  • (21)
@@
  • (22) Kyoto City Office.Kyōto shi tōkeisho, dai jyūsan kai (Statistical Table of Kyoto City, volume. 13). [in Japanese] Kyoto: Kyoto City Office, 1922.
  • (23)
@@
  • (24) Kyoto Prefecture.Kyōto fu tōkeisho, taishō kyū nen, dai san hen (Statistical Table of Kyoto Prefecture, 1920 edition, vol. 3). [in Japanese] Kyoto: Kyoto Prefecture, 1922.
  • (25)
@@
  • (26) Ministry of Agriculture and Commerce. Kōjyō Tsūran (Taishō jyūnen jyūgatsu) (Handbook of Factories, November 1921 edition) [in Japanese]. Tokyo: Ministry of Agriculture and Commerce, 1921.
  • (27)
@@
  • (28) Ministry of Posts and Telecommunications. Denki jigyō yōran, dai jyūyon kai (Handbook of Electricity Industry, vol. 14). [in Japanese] Tokyo: Electric Association, 1922.
  • (29)
@@
  • (30) Nagoya City Office.Nagoya shi tōkeisho, dai nijyūni kai (Statistical Table of Kyoto City, volume. 22). [in Japanese] Nagoya: Nagoya City Office, 1922.
  • (31)
@@
  • (32) Ohkawa, Kazushi, Miyohei Shinohara, and Mataji Umemura. Estimates of Long-term Economic Statistics of Japan since 1868, volume.8. [in Japanese] Tokyo: Toyo Keizai Shinposha, 1967.
  • (33)
@@
  • (34) Osaka City Office. Osakashi tōkeisho, Taishō kyūnen (Statistical Table of Osaka City, 1920 Edition). [in Japanese] Osaka: Osaka City Office, 1922.
  • (35)
@@
  • (36) Postal Savings Bureau. Chochikukyoku nenpō, volume 30 (Annual Statistical Report of the Postal Savings Bureau, volume 30). [in Japanese] Postal Savings Bureau, 1924.
  • (37)
@@
  • (38) Statistics Bureau of the Cabinet. Taishō kyūnen kokusei chōsa hōkoku, zenkoku no bu, dai ikkan (Population Census of Japan, Vol. 1, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1928.
  • (39)
@@
  • (40) Statistics Bureau of the Cabinet. Taoshō kyūnen kokuseichōsahōkoku, zenkoku no bu, dai nikan (Population Census of Japan, Vol. 2, 1920 Edition). [in Japanese] Tokyo: Statistics Bureau of the Cabinet, 1929a.
  • (41)